# All-Electron Molecular Tunnel Ionization Based on the Weak-Field Asymptotic Theory in the Integral Representation

### v1

@@ Appendix A INTEGRAL FORMULAE FOR SINGLE-DETERMINANT WAVE FUNCTIONS
The notation $S_{\sigma}^{\Psi^{+}\Psi}\left[a^{\prime},b^{\prime},\ldots|a,b,\ldots\right]$ denotes the spin-$\sigma$ block of the overlap matrix formed between the orbitals occupied in $|\Psi^{+}\rangle$ removing spin-orbitals $a^{\prime},b^{\prime},\ldots$ and the orbitals occupied in $|\Psi\rangle$ removing spin-orbitals $a,b,\ldots$. The number of orbitals removed from the two wave functions has been ensured such that the resulting modified overlap matrix is square. We will also rewrite the 2-electron potential term in Eq. (20) as
@@ Appendix B OBTAINING OE-WFAT INTEGRAL FORMULA FROM ME-WFAT
Eq. (64) and (B), up to a common constant prefactor, are identical to the integral formula of OE-WFAT in the IR using the occupied orbital $|\psi_{i^{\prime}}^{\sigma}\rangle$ obtained through HF method (compare with eqs. (32)-(34) in [27]). The sign in the prefactor is immaterial, whereas the $N^{-1/2}$ factor will be canceled by the factor $N$ in Eq. (47) upon calculating the ionization rate.

### v2

@@ Appendix A INTEGRAL FORMULAE FOR SINGLE-DETERMINANT WAVE FUNCTIONS
The notation $S_{\sigma}^{\Psi^{+}\Psi}\left[a^{\prime},b^{\prime},\ldots|a,b,\ldots\right]$ denotes the spin-$\sigma$ block of the overlap matrix formed between the orbitals occupied in $|\Psi^{+}\rangle$ removing spin-orbitals $a^{\prime},b^{\prime},\ldots$ and the orbitals occupied in $|\Psi\rangle$ removing spin-orbitals $a,b,\ldots$. The number of orbitals removed from the two wave functions has been ensured such that the resulting modified overlap matrix is square. Note that $\mathcal{R}$ and $\mathcal{S}$ are independent of $\psi_{i}^{\sigma}$, hence $\mathcal{R}$ is a scalar. We will also rewrite the 2-electron potential term in Eq. (20) as
@@ Appendix B OBTAINING OE-WFAT INTEGRAL FORMULA FROM ME-WFAT
Eq. (64) and (B), up to a common constant prefactor, are identical to the integral formula of OE-WFAT in the IR using the occupied orbital $|\psi_{i^{\prime}}^{\sigma}\rangle$ obtained through HF method (compare with eqs. (32)-(34) in Ref. [50]). The sign in the prefactor is immaterial, whereas the $N^{-1/2}$ factor will be canceled by the factor $N$ in Eq. (47) upon calculating the ionization rate.